For the reaction A (aq) + B (aq) <---> C (aq) + D (aq), the equilibrium constant is 24.8 at 25oC and 38.3 at 50oC. What is the value of the change in Gibbs standard free enthalpy (in kJ) of this reaction at 75oC?

1
Expert's answer
2021-11-03T08:13:18-0400

24.8—25

X—75

x=24.8*75/25=74.4


38.3—50

x—75

x=38.3*75/50=57.45


(74.4+57.45)/2=65,925